# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v10

@@ 3.2 Quadratic Gradient Definition
1:training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$, $V\leftarrow\bm{0}$, $W\leftarrow\bm{0}$, $\bar{B}\leftarrow\bm{0}$
4:for $i:=0$ to $d$ do $\triangleright$ $\epsilon$ is a small positive constant such as $81e-$ 8
5:     $\bar{B}[i][i]\leftarrow\epsilon$
6:     for $j:=0$ to $d$ do
7:         $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
8:     end for
9:end for
10:Set $alpha_{0}\leftarrow 0.01$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
11:for $count:=1$ to $\kappa$ do
12:     Set $Z\leftarrow\bm{0}$triangleright $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
13:     for $i:=1$ to $n$ do
14:         for $j:=0$ to $d$ do
15:              $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
16:         end for
17:     end for
18:     Set $\bm{\sigma}\leftarrow\bm{0}$triangleright $\bm{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
19:     for $i:=1$ to $n$ do
20:         $\bm{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
@@
22:     Set $\bm{g}\leftarrow\bm{0}$
23:     for $j:=0$ to $d$ do
@@
25:              $\bm{g}[j]\leftarrow\bm{g}[j]+(1-\bm{\sigma}[i])\times Y[i]\times X[i][j]$
@@
27:     end for
28:     Set $G\leftarrow\bm{0}$
29:     for $j:=0$ to $d$ do
30:         $G[j]\leftarrow\bar{B}[j][j]\times\bm{g}[j]$
@@
32:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$, $\gamma\leftarrow(\frac{10}{1+t})/n$
33:     for $j:=0$ to $d$ do
34:         $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$, $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$, $W[j]\leftarrow w_{temp}$
35:     end for
36:     $alpha_{0}\leftarrow alpha_{1}$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
37:end for
38:Return $V$
@@ 3.3 Quadratic Gradient Algorithms
    $V_{t+1}$ $=\bm{\beta}_{t}+\eta_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
    $\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a). The enhanced NAG is to replace equation 3 with $V_{t+1}=\bm{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually set to $1+\eta_{t}$. Our enhanced NAG is described in Algorithm 1 .
@@
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
@@
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate $N_{t}$ and replaces the gradient with its quadratic form.
@@
(g) The MNIST dataset
(h) The financial dataset
Figure 1: The training results of NAG and Enhanced NAG in the clear.
@@
(g) The MNIST dataset
(h) The financial dataset
Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear.
@@
(g) The MNIST dataset
(h) The financial dataset
Figure 3: The training results of Adam and Enhanced Adam in the clear.
@@
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\bm{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work by Kim et al. (2018a), the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The enhanced Adagrad algorithm utilizes $N_{t}=1+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=1+0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. Refer to (Kingma & Ba, 2014) for further details regarding these parameters. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithm consistently outperforms the original NAG method in terms of convergence rate. Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases. The reason for this discrepancy is beyond the scope of this paper and will be addressed as part of future work. One possible explanation for the superior performance of the enhanced algorithms is that the quadratic gradient integrates curvature information into first-order gradient methods.
@@
An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values. When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge. For quadratic gradient algorithms, we recommend employing an exponentially decaying learning rate, such as $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration number, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive number less than $1$ that controls the rate of decay.
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and the raw quadratic gradient algorithm on the lbw dataset using various learning rates. Figure 4 presents the detailed results of this experiment. It was precisely these results that directly inspired the authors to develop the concept of the quadratic gradient. Similar to the gradient, the quadratic gradient, which incorporates curvature, also demonstrates smooth progression. In particular, as the learning rate is gradually increased, both the gradient and the quadratic gradient reveal corresponding gradual changes in performance, rather than sudden jumps.
@@
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
@@
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
@@
Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset.
@@ 4 Secure Training
We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practical for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
@@
$\text{X}=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$ $\text{ct}_{Z}=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
$\text{ct}_{\beta}^{(0)}=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$ $\text{ct}_{\bar{B}}=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
@@
The public cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
@@
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned shortcomings.
@@
Edin 1253 8 Ours 4 0.04 0.6 89.52 0.943
Baseline 7 0.02 3.6 91.04 0.958
lbw 189 8 Ours 4 0.04 0.6 71.35 0.667
Baseline 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 4 0.31 4.5 79.23 0.637
Baseline 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 4 0.04 0.6 63.20 0.733
Baseline 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 4 0.04 0.6 74.43 0.597
Baseline 7 0.02 3.5 74.44 0.603

### v12

@@ 3.2 Quadratic Gradient Definition
1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
5:  for $i:=0$ to $d$ do
6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
7:     $\bar{B}[i][i]\leftarrow\epsilon$
8:     for $j:=0$ to $d$ do
9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
10:     end for
11:  end for
12:  Set $alpha_{0}\leftarrow 0.01$
13:  Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
14:  for $count:=1$ to $\kappa$ do
15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
16:     Set $Z\leftarrow\boldsymbol{0}$
17:     for $i:=1$ to $n$ do
18:        for $j:=0$ to $d$ do
19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
20:        end for
@@
22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
@@
25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
@@
27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
28:     for $j:=0$ to $d$ do
29:        for $i:=1$ to $n$ do
30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
@@
32:     end for
33:     Set $G\leftarrow\boldsymbol{0}$
34:     for $j:=0$ to $d$ do
35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
36:     end for
37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
38:     Set $\gamma\leftarrow lr/(n\times count)$
39:     for $j:=0$ to $d$ do
40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
42:        $W[j]\leftarrow w_{temp}$
43:     end for
44:     $alpha_{0}\leftarrow alpha_{1}$
45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
46:  end for
47:  Return $V$
@@ 3.3 Quadratic Gradient Algorithms
    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a). The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced NAG algorithm usually set to $1+\eta_{t}$. Our enhanced NAG is described in Algorithm 1 .
@@
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
    $and$
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
@@
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
@@
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 1: The training results of NAG and Enhanced NAG in the clear
@@
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
@@
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 3: The training results of Adam and Enhanced Adam in the clear
@@
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\boldsymbol{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work by Kim et al. (2018a), the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The enhanced Adagrad algorithm utilizes $N_{t}=0.06+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$. Refer to (Kingma & Ba, 2014) for further details regarding these parameters. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
From the empirical results presented in Figures 21, and 3, we conclude that all the enhanced algorithms converge and consistently outperforms the original NAG method in terms of convergence rate.
@@
An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values. When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge. For quadratic gradient algorithms, we recommend employing an exponentially decaying learning rate, such as $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration number, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive number less than $1$ that controls the rate of decay. The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
@@
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
@@
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
@@
Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates. Figure 4 shows the detailed results of this experiment. It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient. Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression. That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
@@ 4 Secure Training
Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
@@
X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
@@
The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
@@
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
@@
Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603